You purchase a 30-year, zero-coupon bond for a price of $35. The bond will pay back $100 after 30 years and make no interim payments. The annual compounded return (geometric average return) on this investment is: OA. 3.0396 OB. 356% OC. 427% OD, 3.38%
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

100=35*(1+r/100)^30

(100/35)^(1/30)=(1+r/100)

(1+r/100)=1.0356

r=1.0356-1

=3.56%(Approx).